# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### v3

@@ 3.1 Supported data types
    None - Sometimes questions are not answerable. With this type it can be checked, whether the model does not answer. Any answer other than None will be penalized.
@@
It is worth mentioning that all combinations of the above types are supported as well. For example, a dictionary may contain lists of strings or the elements of a list may be dictionaries. The implementation of the ANLS* metric maps those complex structures into a tree and compares the ground truth tree against the predicted tree from the model. 0(a) visualizes how the ground truth is decomposed into a tree structure that can then be compared against predictions for an example.
@@
Figure 1: Examples of how the ground truth, as well as predictions, are decomposed into a tree structure. A correct prediction is shown in 0(b), while 0(c) visualizes a partially incorrect prediction. Its worth mentioning that any hallucination as well as incorrect types are penalized as well. More examples are given in Table 1.
@@ 4.2 Experimental Evaluation
In this subsection, the quantitative evaluation of the ANLS* metric on many different datasets and different GLLMs is shown. More precisely, we evaluated two QA datasets (DocVQA [11], MPDocVQA [21]) and five information extraction datasets (Kleister Charity [15], Kleister NDA [15], SROIE [6], VRDU Ad Buy [27], VRDU Registration[27]). The following model versions were used for the evaluation: gpt-3.5-turbo-16k (Version gpt-3.5-turbo-16k-0613), gpt-4-turbo (Version gpt-4-1106-preview) [2], gpt-4-vision (Version gpt-4-vision-preview-1106), gemini-pro (Version 1.0) [17], mistral-large (Feb. 2024) [18] as well as claude-3 (Version claude-3-opus) [16].
We used LangChain [1] for the implementation. Whenever a request to a provider failed (e.g. network timeouts etc.), we send the same request with a sleep time of 10 seconds again. After 5 retries we considered the request as failed resulting in ANLS* of 0.0. We used the following prompts to query each model:
@@
Note that special prompting techniques were required for mistral-large as well as claude-3 in order to get decent values. We refer to the source code for more details. The results are shown in Table 3.
@@
Dataset Method gpt-3.5-turbo-16k gpt-4-turbo gpt-4-vision gemini-pro mistral-large claude-3
DocVQA Simple 0.586 0.607 0.759 0.586 0.445 0.768
Latin Prompting 0.659 0.699 - 0.676 0.447 0.762
SFT (Ours) 0.809 0.790 - 0.741 0.648 0.831
MPDocVQA Simple 0.517 0.635 0.708 0.603 0.364 0.636
Latin Prompting 0.499 0.739 - 0.502 0.335 0.438
SFT (Ours) 0.734 0.781 - 0.616 0.476 0.575
Kleister Charity Simple 0.490 0.743 0.751 0.583 0.652 0.800
Latin Prompting 0.442 0.735 - 0.478 0.576 0.787
SFT (Ours) 0.476 0.763 - 0.633 0.657 0.786
Kleister NDA Simple 0.343 0.695 0.664 0.623 0.637 0.673
Latin Prompting 0.434 0.705 - 0.599 0.624 0.67
SFT (Ours) 0.355 0.703 - 0.552 0.641 0.677
SROIE Simple 0.874 0.835 0.834 0.263 0.855 0.933
Latin Prompting 0.849 0.851 - 0.371 0.863 0.926
SFT (Ours) 0.893 0.873 - 0.288 0.905 0.949
VRDU AD Buy Simple 0.402 0.553 0.640 0.510 0.386 0.577
Latin Prompting 0.389 0.586 - 0.556 0.435 0.608
SFT (Ours) 0.661 0.770 - 0.685 0.594 0.633
VRDU Registration Simple 0.659 0.676 0.665 0.699 0.579 0.685
Latin Prompting 0.693 0.673 - 0.740 0.587 0.715
SFT (Ours) 0.723 0.711 - 0.720 0.639 0.705
@@
First of all, it can be seen that the only model that competes against gpt-4-turbo is claude-3. Both models are good in both, VQA as well as IE tasks. So its not clear yet when to select gpt-4-turbo or claude-3. Nevertheless, it can be seen that gemini-pro, mistral-large as well as gpt-3.5-turbo-16k are quite behind those two models and gpt-4-turbo or claude-3 should be selected for document processing tasks. We believe that Google as well as Mistral will release new competitive versions in the future.
@@
Table Table 3 shows the result for three different prompting methods. The highest scores of 5 out of 7 datasets were achieved with our SFT method. Only for Kleister Charity and Kleister NDA this was not the case. Results for KleisterNDA are very close ($0.703$ vs. $0.705$). For Kleister Charity we found that documents contain mainly text and almost no positional encoding which may describe why SFT and LATIN did not improve simple prompting. Nevertheless, documents usually contain 2D structures and not purely text which shows the importance of specialized prompt formatting techniques such as SFT.
@@
Wang et al. [25] developed and trained a special generative model for documents with 7B parameters. They report 0.634 on DocVQA and 0.499 zero-shot scores on Kleister Charity. Those scores are quite behind our gpt-4-turbo + SFT solution with 0.79 on DocVQA and 0.763 on Kleister Charity. This shows that smaller specialized models are not yet competitive against large purely text-based models that exploit specialized document prompting techniques.

### v4

@@ 3.1 Supported data types
    None - Sometimes questions are not answerable. With this type it can be checked, whether the model does not answer. Any answer other than None will be penalized. On the other hand, if a model generates e.g. a None key in a dictionary that is not in the ground truth, ANLS* ignores it rather than penalizing or rewarding it.
@@
It is worth mentioning that all combinations of the above types are supported as well. For example, a dictionary may contain lists of strings or the elements of a list may be dictionaries. The implementation of the ANLS* metric maps those complex structures into a tree and compares the ground truth tree against the predicted tree from the model. 1(a) visualizes how the ground truth is decomposed into a tree structure that can then be compared against predictions for an example.
@@
Figure 1: Examples of how the ground truth, as well as predictions, are decomposed into a tree structure. A correct prediction is shown in 1(b), while 1(c) visualizes a partially incorrect prediction. Its worth mentioning that any hallucination as well as incorrect types are penalized as well. More examples are given in Table 1.
@@ 4.2 Experimental Evaluation
In this subsection, the quantitative evaluation of the ANLS* metric on many different datasets and different GLLMs is shown. More precisely, we evaluated two QA datasets (DocVQA [11], MPDocVQA [21]) and five information extraction datasets (Kleister Charity [15], Kleister NDA [15], SROIE [6], VRDU Ad Buy [27], VRDU Registration[27]). We used LangChain [1] for the implementation. Whenever a request to a provider failed (e.g. network timeouts etc.), we send the same request with a sleep time of 10 seconds again. After 5 retries we considered the request as failed resulting in ANLS* of 0.0. We used the following prompts to query each model:
@@
Note that special prompting techniques were required for mistral-large as well as claude-3 in order to get decent values. We refer to the source code for more details. The results are shown in Table 5. For gemini-pro we set the safety filter to ”High” to not measure false positives of the safety filter.
@@
Dataset Method gpt-4-turbo gpt-4-vision gemini-1.5-pro mistral-large claude-3
DocVQA Simple 0.589 0.759 0.647 0.445 0.768
Latin Prompting 0.631 - 0.668 0.447 0.762
SFT (Ours) 0.770 - 0.737 0.648 0.831
MPDocVQA Simple 0.606 0.708 0.627 0.364 0.636
Latin Prompting 0.656 - 0.726 0.335 0.438
SFT (Ours) 0.760 - 0.794 0.476 0.575
Kleister Charity Simple 0.708 0.751 0.764 0.652 0.800
Latin Prompting 0.728 - 0.763 0.576 0.787
SFT (Ours) 0.733 - 0.779 0.657 0.786
Kleister NDA Simple 0.633 0.664 0.722 0.637 0.673
Latin Prompting 0.643 - 0.707 0.624 0.670
SFT (Ours) 0.659 - 0.715 0.641 0.677
SROIE Simple 0.849 0.834 0.594 0.855 0.933
Latin Prompting 0.847 - 0.623 0.863 0.926
SFT (Ours) 0.890 - 0.623 0.905 0.949
VRDU AD Buy Simple 0.483 0.640 0.645 0.386 0.577
Latin Prompting 0.506 - 0.686 0.435 0.608
SFT (Ours) 0.712 - 0.765 0.594 0.633
VRDU Registration Simple 0.609 0.665 0.730 0.579 0.685
Latin Prompting 0.614 - 0.749 0.587 0.715
SFT (Ours) 0.590 - 0.780 0.639 0.705
For the evaluation, we always report the values of the newest model version. A detailed comparison between different intermediate versions of GPT as well as Gemini is given in the appendix. The following models were evaluated in this table:
  1. 1.
    gpt-3.5-turbo-16k (gpt-3.5-turbo-16k-0613)
  2. 2.
    gpt-4-turbo (Vgpt-4-1106-preview)
  3. 3.
    gpt-4-vision (gpt-4-vision-preview-1106) [2]
  4. 4.
    gemini-pro (gemini-1.5-pro-preview-0409) [17]
  5. 5.
    mistral-large (Feb. 2024) [18]
  6. 6.
    claude-3 (claude-3-opus) [16]
@@
First of all, it can be seen that the only model that the best performing models are currently claude-3 and gemini-1.5-pro. Both models are good in both, VQA as well as IE tasks. So its not clear yet when to select one over the other. Nevertheless, it can be seen that gpt-4-turbo, gpt-4-vision as well as mistral-large are quite behind those two models. One interesting finding worth mentioning is that gemini-1.5-pro has a quite a restrictive safety filter which explains lower scores on the SROIE dataset as we found that this filter had quite some false-positives on the address field of this dataset. So in future we expect that the same model with an improved safety filter will outperform the other models on even more datasets. A detailed comparison of intermediate versions of GPT as well as Gemini is given in the appendix.
@@
Table 5 shows the result for three different prompting methods. The highest scores of 5 out of 7 datasets were achieved with our SFT method. Only for Kleister Charity and Kleister NDA this was not the case. Results for KleisterNDA are very close. For Kleister Charity we found that documents contain mainly text and almost no positional encoding which may describe why SFT and LATIN did not improve simple prompting. Nevertheless, documents usually contain 2D structures and not purely text which shows the importance of specialized prompt formatting techniques such as SFT.
@@
Wang et al. [25] developed and trained a special generative model for documents with 7B parameters. They report 0.634 on DocVQA and 0.499 zero-shot scores on Kleister Charity. Those scores are quite behind our SFT solution with 0.831 on DocVQA and 0.800 on Kleister Charity. This shows that smaller specialized models are not yet competitive against large purely text-based models that exploit specialized document prompting techniques.
@@ 5 Conclusion & Discussion
[…]
Appendix A Results GPT
In this appendix, a detailed evaluation of different versions of GPT is shown.
Table 4: ANLS* score for different versions of GPT.
Dataset Method gpt-3.5-turbo-16k-0613 gpt-3.5-turbo-0125 gpt-4-1106-preview gpt-4-turbo-2024-04-09 gpt-4-vision-preview-1106
DocVQA Simple 0.576 0.550 0.607 0.589 0.759
Latin Prompting 0.620 0.626 0.699 0.631 -
SFT (Ours) 0.768 0.733 0.790 0.770 -
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708
Latin Prompting 0.489 0.507 0.739 0.656 -
SFT (Ours) 0.724 0.731 0.781 0.760 -
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751
Latin Prompting 0.477 0.493 0.735 0.728 -
SFT (Ours) 0.534 0.548 0.763 0.733 -
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664
Latin Prompting 0.412 0.593 0.705 0.643 -
SFT (Ours) 0.315 0.562 0.703 0.659 -
SROIE Simple 0.509 0.706 0.835 0.849 0.834
Latin Prompting 0.723 0.708 0.851 0.847 -
SFT (Ours) 0.792 0.649 0.873 0.890 -
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640
Latin Prompting 0.424 0.211 0.586 0.506 -
SFT (Ours) 0.676 0.452 0.770 0.712 -
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665
Latin Prompting 0.606 0.630 0.673 0.614 -
SFT (Ours) 0.637 0.654 0.711 0.590 -
Appendix B Results Gemini
In this appendix, a detailed evaluation of different versions of Gemini is shown.
Table 5: ANLS* score for different versions of Gemini.
Dataset Method gemini-1.0-pro-001 gemini-1.5-pro-preview-0409
DocVQA Simple 0.586 0.647
Latin Prompting 0.676 0.668
SFT (Ours) 0.741 0.737
MPDocVQA Simple 0.603 0.627
Latin Prompting 0.502 0.726
SFT (Ours) 0.616 0.794
Kleister Charity Simple 0.583 0.764
Latin Prompting 0.478 0.763
SFT (Ours) 0.633 0.779
Kleister NDA Simple 0.623 0.722
Latin Prompting 0.599 0.707
SFT (Ours) 0.552 0.715
SROIE Simple 0.263 0.594
Latin Prompting 0.371 0.623
SFT (Ours) 0.288 0.623
VRDU AD Buy Simple 0.510 0.645
Latin Prompting 0.556 0.686
SFT (Ours) 0.685 0.765
VRDU Registration Simple 0.699 0.730
Latin Prompting 0.740 0.749
SFT (Ours) 0.720 0.780